OpenStudy (anonymous):

Daniel collected 120 sports cards. Sixty-five of the cards are baseball cards. About what percent of the cards are baseball cards? Round to the nearest percent, if necessary.

jimthompson5910 (jim_thompson5910):

hint: percent of baseball cards = 100*(# of baseball cards)/(# of sports cards)

OpenStudy (anonymous):

huh?

jimthompson5910 (jim_thompson5910):

say you had 50 baseball cards and 200 cards total, that would mean... percent of baseball cards = 100*(# of baseball cards)/(# of sports cards) percent of baseball cards = 100*(50)/(200) percent of baseball cards = 25% Note: this is NOT the answer since this is a similar example (with different numbers)

OpenStudy (anonymous):

forget it ill guess

jimthompson5910 (jim_thompson5910):

just divide the number of baseball cards by the total number of cards then multiply by 100

OpenStudy (anonymous):

ooohhhh ,,, I still don't get it

jimthompson5910 (jim_thompson5910):

what is 65/120 equal to?

OpenStudy (anonymous):

idk

jimthompson5910 (jim_thompson5910):

do you have a calculator?

jimthompson5910 (jim_thompson5910):

if not, then use google https://www.google.com/search?q=65%2F120&oq=65%2F120 don't forget to multiply that result by 100 to get your final answer